OpenStudy (anonymous):

Which of the ordered pairs in the form (x, y) is a solution of this equation? Three x minus begin fraction y over four end fraction equals eleven (3, –8), (4, 4) A. The first is not a solution, but the second is. B. Neither is a solution. C. The first is a solution, but the second is not. D. Both are solutions.

OpenStudy (texaschic101):

3x - y/4 = 11 now we sub in our answer choices to see if they are correct (3,-8)...x = 3 and y = -8 3(3) - 8/4 = 11 9 - 2 = 11 7 = 11....nope...this is incorrect now try (4,4)...x = 4 and y = 4 3(4) - 4/4 = 11 12 - 1 = 11 11 = 11....this one is correct any questions ?

OpenStudy (anonymous):

Thanks! I think I got it!